prove n^3 + 2n is divisible by 3 induction
Answer by ikleyn(52786)

.

The base of induction

    At n= 1  n^3 + 2n = 1^3 + 2*1 = 3  is divisible by 3.

    Thus the base of induction is valid.


The induction step

    Let assume that  P(n) = n^3 + 2n is divisible by 3,



    Then  P(n+1) = (n+1)^3 + 2*(n+1) = n^3 + 3n^2 + 3n + 1 + 2n + 2 = 

                 = (re-group) = (n^3 + 2n) + (3n^2 + 3n + 3) = P(n) + 3*(n^2 + n +1).



    So, P(n+1) is the sum of P(N) and the other addend, which is multiple of 3.


    Thus, if P(n) is divisible by n, then P(n+1) is divisible by 3, too.


    The inductive step is proven to be valid.


Hence, according to the Mathematical induction principle, the statement is true for all positive integer n.

The proof is completed.

----------------------

On the method of Mathematical induction see the lessons
    - Mathematical induction and arithmetic progressions
    - Mathematical induction and geometric progressions
    - Mathematical induction for sequences other than arithmetic or geometric
    - Proving inequalities by the method of Mathematical Induction
    - OVERVIEW of lessons on the Method of Mathematical induction

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic
"Method of Mathematical induction".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.